A square ammeter has sides of length 3.00 cm. The
Answer #1

4. Applying Ampere´s law,

\oint \vec{B}.d\vec{l}= \mu _{0}I_____________________________(1)

we choose a square path,

B4l= \mu _{0}I,\, \, \, \, \, I= \frac{4Bl}{\mu _{0}}_____________________(2)

using (2), the current is

I= \frac{4\times 4\times 10^{-4}\, T\times 3\times 10^{-2}\, m}{4\pi \times 10^{-7}\frac{Tm}{A}}=0.38 \, A

______________________________________________

5) The magnetic field is given by

\vec{B}= \vec{B}_{1}+\vec{B}_{2},_____________________________(1)

where,

\oint \vec{B}_{j}.d\vec{l}= \mu _{0}I_{j},\, \, \, \, \, \, j= 1,2_____________________(2)

The magnetic field inside is

\oint \vec{B}_{1}.d\vec{l}= \mu _{0}I_{1},\, \, \, \, \, \, \, \vec{B}_{1}= \frac{\mu _{0}I_{1}}{2\pi r}\hat{\theta }_________________(3)

\oint \vec{B}_{2}.d\vec{l}= \mu _{0}I_{2},\, \, \, \, \, \, \, \vec{B}_{2}= \frac{\mu _{0}I_{2}}{2\pi r}\left ( -\hat{\theta } \right )______________(4)

substituting (3) and (4) into (1), yields (inside)

\vec{B}_{in}= \frac{\mu _{0}}{2\pi r_{i}} \left ( I_{1}-I_{2} \right ) \hat{\theta }= \frac{4\pi\times 10^{-7} \times 2.5}{2\pi \times 0.5\times 10^{-2}\, }\, T\, \hat{\theta }= 10^{-4}\, T\, \hat{\theta }____________________(5)

_________________________

The magnetic field outside the ciindrical conductor

\vec{B}_{out}=\left (\frac{\mu _{0}I_{1}}{2\pi r_{1}}-\frac{\mu _{0}I_{2}}{2\pi r_{2}} \right ) \hat{\theta }= -0.66\times 10^{-5}\, T\, \hat{\theta }

___________________________________________________